OpenStudy (logan13):

What is the area of a circle with a radius of 5 inches? (Use = 3.14.) 31.4 square inches 78.5 square inches 15.7 square inches 314 square inches

OpenStudy (logan13):

https://app159.studyisland.com/images/blank.gif

OpenStudy (unanimoose):

You are so ridiculous.

OpenStudy (logan13):

@AlexandervonHumboldt2

OpenStudy (anonymous):

b

OpenStudy (unanimoose):

No matter how much we try to teach you how to do it by yourself you just keep posting them here for the answer.

OpenStudy (anonymous):

Its b fan and medal

OpenStudy (anonymous):

Area = pi * r^2 Area = 3.14 * (5^2) Area = 3.14 * 25 = 78.5 B, 78.5 square inches

OpenStudy (unanimoose):

Try and learn something from what we teach you so that you don't have to post everything here.

OpenStudy (unanimoose):

And @IQ250 , please don't give a direct answer.

OpenStudy (logan13):

@AlexandervonHumboldt2

OpenStudy (anonymous):

lol

OpenStudy (alexandervonhumboldt2):

i hope you undestood all